# GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness

### v3

@@ 4.2.1 Layer-Channel Scaling
[…]
4.2.2 Vectorized Quantization
In the layer-channel scaling stage, the activation tensor of each layer is vectorized along the channel dimension. In the subsequent vectorized quantization stage, both the scaling factor $\vec{s}$ and zero-point $\vec{z}$ are also vectorized and computed as shown in Equation 5. The activation values go through quantization and dequantization (Equation 6), with all operations running in parallel across channels. During computation, scaling factor $s$ and zero-point $z$ of each channel are broadcasted to the activation values and applied element-wise.
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$ (5)
$x_{q,l}=\left\lfloor\frac{x}{\vec{s}}+\vec{z}\right\rceil,x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$ (6)
This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors. A detailed computational process is provided in Algorithm 1, available in the supplementary material.
5 Experiments
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets. For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet. All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26]. For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35]. Further implementation details are provided in the supplementary material, Section B. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@ 5.1 Experimental Setup and Details
Layer-Q Channel-Q
Layer-Channel
Layer-Channel
(Vectorization, GranQ)
Accuracy (%)
49.98 (±$\pm 0.24$)
51.26 (±$\pm 0.28$)
62.68 (±$\pm 0.11$)
62.73 (±$\pm 0.17$)
Table 5: Comparative analysis of four quantization approaches on ResNet-20 accuracy. The experiments were conducted using the AdaDFQ method on CIFAR-100 with 3-bit quantization. Results are averaged over three repetitions with standard deviation.
@@ 5.2 Performance Evaluation
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [22] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [41]. In contrast, GranQ maintains the speed advantage of GAN-based generation while achieving high performance through fine-grained adjustments.
@@ 5.3.1 Effectiveness Evaluation
Figure 5: Ablation study on the latency of the quantization process in ResNet-20 across different batch sizes. The experiments were conducted using the AdaDFQ method on CIFAR-100 with a 3-bit quantization setting.
@@ 5.3.2 Efficiency Evaluation
In general, performing precise quantization that considers both layers and channels leads to an exponential increase in computational cost. However, GranQ is designed to minimize this overhead while maintaining high accuracy by leveraging vectorized operations. To validate this, we measured the latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares four methods.
The first two methods are conventional quantization techniques: layer- and channel-wise. The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations. The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization. Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3. In contrast, GranQ enhances efficiency through parallelized computations (Algorithm 1 in supplementary material).
A key observation from Figure 5 is that GranQ induces only a slight increase in latency compared with layer- and channel-wise quantization. Moreover, as summarized in Table 5, GranQ achieves significantly higher accuracy than traditional quantization methods. Even when compared with the less efficient iteration-based layer-channel quantization, GranQ maintains nearly identical accuracy. Additionally, as summarized in Table 2, GranQ substantially reduces quantization error and effectively preserves activations. These findings confirm that GranQ achieves fine-grained quantization while benefiting from parallel computation.
@@ 6.1 Why GranQ is Effective?
The most common activation quantization approaches are known for their low computational cost and fast processing speeds. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 3). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section D in the supplementary material). These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
To overcome these issues, GranQ introduces the layer-channel scaling method, which combines layer- and channel-wise quantization. Unlike conventional methods, GranQ considers the activation distribution of each channel individually while also incorporating layer information, enabling more precise quantization. Furthermore, although conventional quantization relies on scalar operations with a single scaling factor, GranQ utilizes multiple scaling factors and performs vectorized computations. This approach leverages GPU resources efficiently, minimizing computational overhead while achieving effective quantization.
Experimental results demonstrate that GranQ outperforms traditional layer- and channel-wise quantization methods, achieving higher cosine similarity and lower relative error (Table 2). This demonstrates that GranQ accurately preserves the original activation distribution. Additionally, GranQ maintains computational efficiency in terms of latency and effectively balances the trade-off between computational cost and performance (Figure 5 and Table 5). Most notably, GranQ achieves SOTA performance compared with those of existing layer-wise quantization methods (Table 3), highlighting its efficiency and superiority over scalar iteration approaches.
@@ 6.2 Limitation and Future-work
As GranQ aims to precisely preserve activations during quantization, it partially depends on generated data quality. Consequently, studies on data generation remain highly important. The proposed method has the potential to achieve even better performance when applied to high-quality data generation strategies. In future work, we aim to improve the quantization performance by focusing on both the accurate generation of original activations and their effective preservation.
@@ 7 Conclusion
We present a comprehensive analysis of activation quantization errors in ZSQ and propose GranQ, a novel quantization method to effectively reduce these errors. Conventional layer- and channel-wise quantization methods rely on a single scaling factor, which often results in substantial activation loss. In contrast, GranQ employs a fine-grained quantization approach that simultaneously considers both layer and channel information, thereby reducing activation loss and enabling more precise quantization. Moreover, GranQ successfully maintains computational efficiency through vectorized quantization while performing detailed and fine-grained quantization.
@@
Acknowledgement This research was supported by the National Research Foundation (NRF) funded by the Korean government (MSIT) (No. RS-2023-00229822).
Supplementary Material
Appendix A Notation
@@
$x$ Original activation value
$x_{q}$ Quantized activation value
$x_{\text{deq}}$ Dequantized activation value
$x_{\min}$, $x_{\max}$ Minimum and maximum activation values
@@
$\min\limits_{h,w}(x)$, $\max\limits_{h,w}(x)$ Channel-wise min/max activations
@@
In this section, we provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($x$), quantized activation ($x_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
Appendix B Details of Experimental Setup
We provide a detailed description of the datasets, experimental setup, and comparison methods utilized in our experiments. This ensures the implementation of the experiments and clearly highlights the strengths of the proposed approach over existing methods.
B.1 Datasets
@@
B.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [14]. Specifically, ACGAN [12] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [14]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [11] as the optimizer, with a momentum of 0.9 and a weight decay of 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [13], and visualizations were generated with matplotlib [7]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
B.3 Comparison Methods
@@
Algorithm 1 GranQ: Layer-Channel Quantization
1:Activation tensors $\{x_{l}\}_{l=1}^{L}$, where $x_{l}\in\mathbb{R}^{C\times(HW)}$, quantization bit $b$
2:Dequantized activation tensors $\{x_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do $\triangleright$ Process each layer separately
4:       Compute per-channel min/max values:
$\vec{x}_{\min}=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$
$\triangleright$ Compute along the channel axis for layer $l$
5:       Compute scaling factors and zero-points (Vectorized):
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$
6:       Quantize values (Vectorized):
$x_{q,l}=\left\lfloor\frac{x_{l}}{\vec{s}}+\vec{z}\right\rceil$
$\triangleright$ Element-wise quantization for all channels in layer $l$
7:       Dequantize values (Vectorized):
$x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$
$\triangleright$ Reconstruct activation tensor for layer $l$
8:end for
9:Return the dequantized tensors:
$\textbf{return}\ \{x_{\text{deq},l}\}_{l=1}^{L}$
@@
  • PLF [5]: PLF was the first to evaluate generated data before quantization, offering a method to ensure more reliable data.
@@
Figure 6: Ablation study on relative error in ResNet-20 across layers. Layer-Q, Channel-Q, and GranQ represent layer-wise, channel-wise, and proposed quantization. Experiments used AdaDFQ on CIFAR-100 with 3-bit quantization.
Appendix C Algorithm
GranQ is a granular quantization method that processes activation values across all channels within each layer. Algorithm 1 describes the procedure for quantizing the activation tensor $\{x_{deq,l}\}^{L}_{l=1}$. The activation tensor of layer $l$ has a shape of $C\times(HW)$, and the minimum and maximum values are calculated independently for each channel. Based on these values, a scaling factor and zero-point are computed per channel for the quantization process. The resulting quantized values are dequantized back to floating-point representation prior to returning.
Appendix D Extra Analysis
D.1 Analysis of Activation Preservation
To analyze the severe quantization error (activation distortion) occurring in specific layers, we conducted additional experiments, as illustrated in Figure 6. This figure presents the relative error for each layer of ResNet-20 and reveals that certain layers such as the 18th exhibit significantly high errors. These findings suggest that both layer- and channel-wise quantization fail to effectively preserve activations in specific layers. In contrast, GranQ achieves consistently low errors across all layers, ensuring better activation preservation.
D.2 Analysis of Activation Magnitude
In Figure 7, we observe that channel-wise quantization is ineffective in ZSQ settings, similar to layer-wise quantization. As shown in Figure 7(a), the full-precision (FP) model exhibits significant activation magnitude differences across layers, even within the same channel axis. This highlights the limitations of channel-wise quantization in handling quantization errors. Conversely, GranQ addresses this issue by accounting for each channel across all layers, thereby minimizing quantization errors. This is particularly important in the final layers, where maintaining the activations of the FP model is critical for accurate inference. Whereas (a) fails to preserve these distributions, (b) successfully maintains them.
(a) Layer Quantization

### v4

@@ 5.1 Experimental Setup and Details
[…]
Table 4: Accuracy of existing SOTA methods with the integration of GranQ. GDFQ and AdaDFQ focus on data generation, whereas Qimera+AIT and AdaDFQ+AKT primarily enhance quantized model training.
Figure 4: Latency of ResNet-20 quantization across batch sizes on CIFAR-100 with 3-bit setting.
Figure 5: Relative quantization error across layers in ResNet-20 with 3-bit quantization on CIFAR-100.
@@ 5.2 Performance Evaluation
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [46] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [16]. In contrast, GranQ improves the quantization mechanism without relying heavily on diffusion-based data generation, as GenQ does. As our method is adaptable to existing QAT frameworks, it can potentially achieve even higher performance when combined with GenQ’s data generation.
@@ 5.3.2 Efficiency Evaluation
Method
Quantization
Application
Scaling Factor
Computation
Accuracy (%)
Quantization
Latency (ms)
Layer-wise 49.98 ($\pm$ 0.24) 0.227
Channel-wise (Conventional scalar-loop) 62.68 ($\pm$ 0.11) 103.671
Channel-wise (Vectorized, GranQ) 62.73 ($\pm$ 0.17) 0.236
Table 5: Ablation study of vectorization in scaling computation and quantization. GranQ applies vectorization to both stages, achieving high accuracy with low latency. Latency is measured with batch size 16. Experiments use AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation. By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations. To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels (Algorithm 1 in appendix).
The two channel-wise methods perform per-channel quantization. The conventional method computes scaling factors sequentially, whereas GranQ vectorizes the scaling computation across channels for improved efficiency. A key observation from Figure 5 is that GranQ incurs only a slight increase in latency compared to conventional layer-wise quantization. With only a minor increase in latency, GranQ achieves higher accuracy than the layer-wise baseline, as shown in Table 5. Compared to scalar-loop-based channel quantization, which implements the same computation sequentially, GranQ achieves nearly identical accuracy. Furthermore, Table 2 shows that GranQ substantially reduces quantization error and more effectively preserves activation information. These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency. Additional analysis and experimental results on inference efficiency are provided in Table 7 of the appendix.
@@ 6.1 Why GranQ is Effective?
The most common approaches to activation quantization (i.e., layer-wise quantization) are known for their low computational cost and fast processing speed. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 2). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts.
To overcome these issues, GranQ computes and stores per-channel scaling factors for each input in a unified vector form, enabling parallel processing across all channels. Unlike conventional methods, it performs quantization based on each input’s channel-wise distribution, resulting in more precise and efficient computation.
@@ 6.2 Limitation and Future-work
In ZSQ settings, synthetic data often exhibit broader and more distorted activation distributions compared to original data (Figure 6 in appendix). GranQ addresses this challenge by applying dynamic quantization for each channel based on its activation distribution, making it particularly effective in QAT scenarios. However, GranQ is less suitable for PTQ, which relies on static activation quantization. As future work, we plan to develop quantization methods that preserve the fine-grained scaling of GranQ while enhancing adaptability to PTQ settings.
@@ 7 Conclusion
Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training. In contrast, channel-wise scaling in activation quantization is more challenging, as activation values vary with each input, making it difficult to apply fine-grained scaling. These limitations become more pronounced in QAT settings, where large amounts of data are used. Under these conditions, quantization is strongly affected by activation distributions. This issue is especially critical in ZSQ, where training data are synthesized using internal statistics from the FP model. Synthetic data generated in this way often produce distorted activation distributions compared to original data, resulting in increased quantization error during scaling. To address these issues, we propose GranQ, a dynamic channel-wise quantization method that adapts scaling factors to each input’s activation distribution. This enables more precise quantization even under distorted distributions, as often observed in ZSQ. In addition, GranQ improves computational efficiency by applying per-input, per-channel scaling through vectorized operations, which is rarely adopted in activation quantization due to its implementation complexity.
@@
Appendix A Appendix
A.1 Notation
We provide an overview of the main symbols and definitions employed throughout the paper. It specifically includes essential elements such as the original activation ($A$), quantized activation ($A_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
@@ A.1 Notation
$A$ Original activation value
$A_{q}$ Quantized activation value
$A_{\text{deq}}$ Dequantized activation value
$A_{\min}$, $A_{\max}$ Minimum and maximum activation values
@@
$\min\limits_{h,w}(A)$, $\max\limits_{h,w}(A)$ Channel-wise min/max activations
@@
A.2 Algorithm
While prior works apply vectorized computation during the quantization step, they typically compute per-channel scaling parameters using iterative channel-wise operations. GranQ differs by also vectorizing the scaling computation, allowing both steps to be processed in parallel. Algorithm 1 outlines the complete process of quantizing and dequantizing activation tensors using GranQ.
In Algorithm 1, both the scaling and quantization steps are implemented using vectorized tensor operations across all channels. Specifically, the per-channel minimum and maximum values are computed simultaneously, and the resulting scaling factors and zero-points are applied through broadcasting. This eliminates the need for channel-wise loops and allows efficient execution on modern hardware accelerators.
In contrast, Algorithm 2 follows a conventional channel-wise quantization approach, where the scaling parameters are computed sequentially for each channel using an explicit for-loop. The quantization and dequantization steps are still vectorized per channel. However, the scaling computation remains scalar and iterative. This leads to increased computational overhead, especially when applied to deep networks with many channels. By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
Algorithm 1 GranQ: Vectorized Per-Channel Scaling for Activation Quantization
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:    Reshape:
$A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
$\triangleright$ Flatten spatial dimensions
5:    Decompose by channel:
$\vec{A}_{\min}=\min_{h,w}(A_{l}),\quad\vec{A}_{\max}=\max_{h,w}(A_{l})$
$\triangleright$ Extract per-channel range
6:    Compute scaling parameters:
$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$
7:    Quantize (vectorized):
$A_{q,l}=\left\lfloor\frac{A_{l}}{\vec{s}}+\vec{z}\right\rceil$
8:    Dequantize (vectorized):
$A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$
9:end for
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
Algorithm 2 Conventional Channel-Wise Activation Quantization (Scalar-loop for scaling)
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:    Compute per-channel scaling parameters:
5:    for each channel $c$ in $C$ do
$s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1},\quad z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
6:    end for
7:    Apply quantization (vectorized):
$A_{q,l}[c,:,:]=\left\lfloor\frac{A_{l}[c,:,:]}{s_{c}}+z_{c}\right\rceil,\quad\forall c$
8:    Dequantize (vectorized):
$A_{\text{deq},l}[c,:,:]=s_{c}(A_{q,l}[c,:,:]-z_{c}),\quad\forall c$
9:end for
10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
A.3 Details of Experimental Setup
We describe the datasets, experimental setup, and baselines to ensure reproducibility and highlight both the effectiveness and implementation accessibility of our approach.
A.3.1 Datasets
@@
A.3.2 Experimental Setup
The data generation process followed the settings of AdaDFQ [1]. Specifically, ACGAN [2] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3. All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [1]. For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet. The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100. All experiments used Nesterov-based SGD [47] as the optimizer, with momentum 0.9 and weight decay 1e-4. In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs). GranQ was implemented using PyTorch [3], and visualizations were generated with matplotlib [4]. All experiments were conducted on an NVIDIA RTX 3090 GPU.
A.3.3 Comparison Methods
@@
These methods cover various aspects of ZSQ, including data generation, training objectives. We compare them to demonstrate the effectiveness and generality of GranQ, which focuses on improving the quantization process itself by introducing an optimized, vectorized scaling mechanism. This structural enhancement allows GranQ to be seamlessly integrated into existing QAT-based ZSQ pipelines, providing consistent gains across different settings. Furthermore, as several methods have not yet released official code, we expect that GranQ can be readily applied as their implementations become publicly available.
A.4 Extra Analysis
A.4.1 Data Distribution Analysis
(a) Original inputs (Skewness: 0.2599)
(b) Synthetic inputs (Skewness: 0.4877)
Figure 6: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet. The synthetic inputs, generated by AdaDFQ [17], exhibit higher skewness and a rightward shift, indicating the presence of more extreme activation values and motivating the need for granular scaling in quantization.
Figure 6 compares the pixel distributions of original and synthetic inputs derived from ImageNet. The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values. This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values. As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle. In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion. These observations underscore the limitations of conventional scaling strategies. To address this, we propose a granular, channel-wise scaling approach that dynamically adapts to per-channel statistics and better preserves activation fidelity under distributional mismatch.
A.4.2 Computational Efficiency Analysis
Although GranQ and conventional channel-wise quantization share a higher theoretical complexity of $\mathcal{O}(L\cdot C\cdot H\cdot W)$ compared to the layer-wise baseline with $\mathcal{O}(L\cdot H\cdot W)$, all three methods apply their respective scaling strategies only during the training phase. Since QAT fixes the scaling factors as static parameters prior to deployment, these operations are not executed during inference, resulting in no runtime overhead. Consequently, the per-channel granularity used in GranQ does not incur any additional inference latency. As shown in Table 7, all methods achieve comparable inference latency regardless of the scaling strategy, demonstrating that GranQ maintains runtime efficiency while benefiting from fine-grained scaling during training.
Table 7: Comparison of activation quantization latency on ResNet-18 (ImageNet, batch size = 128) under a 3-bit quantization setting. All latency values are measured per batch and averaged over 3 evaluations, with standard deviation reported in parentheses. Theoretical complexity reflects quantization-time operations and does not affect inference latency, which is measured separately.
Method Theoretical Complexity Scaling Strategy Inference Latency (ms)
Activation (Layer-wise) $O(L\cdot H\cdot W)$ Layer-wise uniform scaling 44.194 (+0.190)
Activation (Channel-wise, Scalar-loop) $O(L\cdot C\cdot H\cdot W)$ Channel-wise scalar scaling (loop-based) 44.272 (+0.158)
GranQ (Channel-wise, Vectorization) $O(L\cdot C\cdot H\cdot W)$ Channel-wise vectorized scaling 44.231 (+0.186)
A.4.3 Analysis of Activation Preservation
We analyzed how activation distortion propagates across layers in a ZSQ setting where synthetic data is used as input. When coarse quantization is applied without sufficiently granular scaling, distortion tends to accumulate in specific layers and ultimately degrades the output layer’s activation representation. This suggests that insufficient preservation of intermediate activations can negatively affect the quality of final inference. The problem becomes more severe in zero-shot scenarios due to the distributional mismatch introduced by synthetic data. In contrast, GranQ mitigates this issue by applying fine-grained channel-wise scaling, resulting in consistently stable activation distributions throughout the network. This phenomenon is clearly demonstrated in Figure 7, where GranQ maintains stable activations with noticeably reduced distortion compared to conventional quantization methods.
(a) Layer-wise Quantization